OpenStudy (anonymous):

2x − 2y − 5z + 2x + 2y + 3z I need help simplify

OpenStudy (marissalovescats):

Combining like terms!

OpenStudy (marissalovescats):

2x+2x=? -2y+2y=? -5x+3z=?

OpenStudy (anonymous):

2x+2x=4x -2y+2y=0 -5x+3z=-2 its tha correct

OpenStudy (marissalovescats):

Yup! So your answer is 4x-2z

OpenStudy (anonymous):

truth